OpenStudy (anonymous):

sqrt{2x}=2+sqrt{x-2}

OpenStudy (wasiqss):

plz man do sumthing by urself!

OpenStudy (diyadiya):

\[\sqrt{2x}=2+\sqrt{x-2}\]

OpenStudy (diyadiya):

Square Both the Sides

OpenStudy (anonymous):

sqrt{2x}-sqrt{x-2}=2 2x+x-2+2sqrt{2x^2-4x}=4 2sqrt{2x^2-4x}=3x-6 x^2-2x+36=0 x1=18 x2=2